# Chatbot Companionship: A Mixed-Methods Study of Companion Chatbot Usage Patterns and Their Relationship to Loneliness in Active Users

### v1

@@ 3.1.2 Custom Survey Items
In addition to established scales, we developed custom items to capture specific aspects of chatbot usage:
  • Usage Characteristics: Frequency of chatbot use and average session length.
  • Motivations and Continued Use: Multiple-choice questions with required free-text elaboration on why participants began using chatbots and why they continue to use them.
  • Human vs. Chatbot Preference: Several items exploring participants’ preferences for human or chatbot interaction in various contexts.
  • Conversation Topics: Questions about the topics participants discuss with chatbots.
  • Perceived Impact on Relationships: Likert-scale items assessing participants’ perceptions of how chatbot use has affected their relationships with people.
These custom items were designed to provide context-specific data on chatbot usage patterns and perceived effects, complementing the insights gained from established scales.
@@ 3.2 Correlation and Multiple Regression
Data processing and statistical analyses were performed using Python, with the statsmodels library for regression analyses and scikit-learn for data preprocessing. We used a significance level of $\alpha=0.05$ for all statistical tests, unless otherwise specified. Respondents’ frequency of chatbot usage and session length were converted to linear numerical scores; any multiple responses were averaged. Numerical variables, including the frequency and session length score, were scaled to have a mean of 0 and a standard deviation of 1 using StandardScaler to improve interpretability and reduce potential multicollinearity [Aiken and West(1991)]. We focused on chatbot session length as the independent variable representing chatbot usage in the model, as it had a significant correlation with loneliness in the correlation analysis where frequency did not. We computed a Spearman correlation matrix for all standardized variables and calculated corresponding p-values to identify significant relationships (p < 0.05) between variables of interest [Cohen et al.(2003)]. The p-values were then corrected using the Benjamini-Hochberg procedure [Benjamini and Hochberg(1995)]. For the following analyses, we focused on loneliness (ULS-8) as the dependent variable and session length score as the primary independent variable, though frequency score was also included in our process of model exploration.
Potential mediators were identified using the corrected p-values (p < 0.05). We conducted mediation analysis following Baron and Kenny’s [Baron and Kenny(1986)] approach, regressing the dependent variable on the independent variable, the mediator on the independent variable, and the dependent variable on both the independent variable and mediator. We calculated indirect effects and conducted Sobel tests [Sobel(1982)] to assess the significance of mediation (p < 0.05).
We identified potential moderators using the corrected p-values (p < 0.05). To assess moderation effects, we conducted moderation analyses following the approach outlined by Aiken and West [Aiken and West(1991)]. For each potential moderator, we created an interaction term with the independent variable and fitted a linear regression model including the independent variable, moderator, and their interaction. The significance of the interaction term (p < 0.05) was used to determine the presence of moderation effects.
To account for potential confounding variables, we implemented a data-driven approach to identify and control for confounders [Greenland et al.(1999), VanderWeele and Shpitser(2013)]. We examined the correlation matrix and corresponding p-values to identify variables that were significantly correlated (corrected $p<0.05$) with both the independent and dependent variables.
@@
To address potential multicollinearity, we calculated Variance Inflation Factors (VIF), considering variables with VIF > 5 for elimination [Hair et al.(2010), O’brien(2007)]. After removing the state empathy subscales, which were redundant with the full scale, all predictors were below the threshold.
@@
We explored various regression models using Ordinary Least Squares (OLS) regression, beginning with all identified mediators, moderators, and confounders. We included socialization (LSNS) as a potential mediator and neuroticism as a potential confounder on a theoretical basis, as socialization is the objective counterpart to loneliness, and neuroticism is strongly linked to loneliness [Buecker et al.(2020)] and may act as a driver for emotional regulation using chatbots. While retaining age and sex as control variables, we iteratively refined the model by removing predictors with high p-values (generally p > 0.1) in a process of backward elimination, to make a more parsimonious and interpretable model [Cohen et al.(2003)]. After elimination, we experimented with adding social attraction to a close person, and its interaction with session length, to the model. For each of the remaining predictors, we explored adding one new interaction term at a time, with significant interaction becoming part of the model. Frequency score was removed from the model, as its p-value (0.636) was far above the significant threshold. Categorical variables (area type - rural/urban/suburban, education, relationship/marital status, political party, race, employment status) were tested using ANOVA one at a time but were not retained in the final model due to non-significance.
@@
To examine the potential mediators of our final model, we conducted a parallel multiple mediation analysis. This analysis was performed using a bootstrapping approach with 5,000 resamples, as recommended by [Preacher and Hayes(2008)]. The analysis controlled for several covariates, including age, sex, and personality traits (BFI). Confidence intervals (95%) for the indirect effects were calculated using the percentile method. All analyses were conducted using Python with the statsmodels library.
@@ 3.3 Cluster Analysis
To identify distinct user profiles of chatbot users based on their psychosocial characteristics, we employed K-means clustering analysis. These profiles were identified based on a set of features with significant correlations (p > 0.05, corr > 0.3) to key variables, as well as features relevant to chatbot usage including usage frequency and measures of empathy and attraction with a chatbot; the features were reduced to 10 principal components. After some exploration, we used Chatbot Session Length, Frequency, Loneliness (ULS-8), and Problematic Use of Chatbot (GPIUS-2) for our key variables. The optimal number of clusters was determined using a combination of quantitative metrics and qualitative interpretation:
@@
The final clustering was performed using the K-means algorithm with k=7, implemented in Python using the scikit-learn library [Pedregosa et al.(2012)]. The resulting clusters were then characterized based on their centroids and distribution of key variables.
To validate our clustering solution, we employed a multi-faceted approach:
@@ 3.4 Interpretation of Quantitative Results and Free-Text Responses
Second, we examined the free-text responses associated with multiple-choice items to complement our quantitative findings. While not a formal qualitative analysis, this examination allowed us to contextualize our quantitative results and provide illustrative examples of user experiences within each cluster and across the entire sample.
@@ 4 Results
4.1 Participant Demographics and Usage Characteristics
We surveyed 404 companion chatbot users recruited via CloudResearch. Participants ranged in age from 18 to 73, with the distribution shown in Figure 5, and were primarily from the United States. Additional demographic characteristics are visualized in Figure 4.
In addition to the participants’ demographics, we also collected information about their usage, including which companion chatbots they use, how much they use the chatbots, why they use the chatbots, and what they use the chatbots for. Most participants utilize Snapchat MyAI (31.63%), Character.ai (30.39%), or Replika (22.61%) (see Figure 6). Most respondents reported spending 15 and 30 minutes in each session with a companion chatbot (42.64%), with 35.39% reported sessions shorter than 15 minutes. A smaller portion of users engage for 30 minutes to an hour (15.14%), and only 4.90% of participants reported sessions lasting between 1 and 2 hours. Very few users spend extended periods chatting, with only 1.28% of sessions exceeding 4 hours, and 0.64% lasting between 2 and 4 hours. Additional characteristics can be seen in Figure 6 and Figure 7.
Figure 4: Comprehensive demographic breakdown of study participants. Six pie charts illustrate the distribution of gender, race, relationship/marital status, household income, country of residence, and employment status.
Figure 5: Participants’ age distribution. Participants were 18 to 73 years of age.
4.1.1 Motivations for initial interest and continued use of companion chatbots.
The reasons participants use and continue to engage with companion chatbots are diverse (see Figure 7), with non-social motivations predominating. Curiosity about AI and chatbots (30.89% initial use, 20.91% continued use) is a primary driver. As one user stated, “I was curious to see what they were like,” capturing the initial intrigue many feel. Others sought to explore the technology’s capabilities: “I wanted to see if it could actually emulate a real person.”
Figure 6: Overview of chatbot usage patterns among participants. The graph on the left shows the most used chatbot platforms, with 31.6% of participants using Snapchat MyAI, 30.4% using Character.ai, and 22.6% using Replika, while smaller portions use other platforms such as Anima (4.8%) and Xiaoice (1.2%). The center graph presents the chatbot session length, where 42.6% of participants use chatbots for 15–30 minutes, 35.4% for less than 15 minutes, and 15.1% for 30 minutes to 1 hour. The right graph displays chatbot usage frequency, with 43.9% using chatbots for a few months, 35.2% using them for several months to around a year, and smaller percentages for longer durations such as multiple years (7.7%).
Figure 7: Analysis of participant responses regarding their use of companion chatbots. The top-left graph presents reasons for using companion chatbots, where 30.9% of participants expressed curiosity about AI and chatbots, followed by 26.4% who sought fun and entertainment. The top-right graph displays conversation topics with chatbots, with casual conversation being the most common (26.3%), followed by entertainment and play (21.7%). The bottom-left graph shows motivations for continued use, where enjoyment/fun was the leading factor (23.7%), followed by satisfying curiosity (20.9%). Lastly, the bottom-right graph illustrates preferred devices for companion chatbot use, with the majority of users favoring smartphones (36.9%), followed by computers (30.9%).
Fun and entertainment (26.35% initial, 23.69% continued) are significant factors. Users enjoy role-playing with characters and find humor in interactions: “I’ve watched YouTubers use Character.ai before and found it to be a humorous experience.” The entertainment often intertwines with creativity (17.81% initial use), with users employing chatbots for storytelling and game preparation: “I use chatbots to practice conversations for my D&D game, creating NPC personalities and dialogue.”
Passing time and reducing stress (12.88% continued use) are notable reasons for ongoing engagement. Users find chatbots helpful for relaxation: “When I’m stressed or bored, chatting with an AI feels like a safe and fun way to distract myself.”
Social motivations, while less prevalent, are significant for some users. These include coping with loneliness (12.24% initial, 6.92% continued), improving social and conversational skills (7.24% initial), and seeking companionship (9.86% continued). One user noted, “I started using Replika during the COVID-19 pandemic when I was lonely and seeking some form of social interaction.” For others, chatbots offer a low-pressure environment to practice social skills: “I’m autistic, so I’m not the best at social interaction, and I found that practicing with chatbots helped.”
A small portion of users (3.53% initial) seek relationship-like interactions with chatbots, exploring emotional connections in a safe space. Continued use is also motivated by emotional experiences (6.52%), self-discovery (6.12%), understanding others (5.88%), and boosting self-confidence (5.41%). One user shared, “I’ve learned a lot about myself from talking to the chatbot – it helps me think through things in a different way;” another noted that “Chatbots give me different perspectives that I wouldn’t have considered on my own, which helps me when interacting with others in real life,” demonstrating how the social simulation offered by companion chatbots offer the opportunity to improve both interpersonal and intrapersonal understanding.
@@ 4.1.2 Topics of conversation with companion chatbots
The range of topics discussed with chatbots is diverse, reflecting users’ varied interests and needs. Casual conversations (26.28%) dominate, with users engaging in everyday chat without social pressure: “I often chat with my AI about random things like my day, and it feels like I’m having a normal conversation without the pressure of talking to a real person.”
Entertainment and play (21.66%) are common, especially for users interested in role-playing or creative writing: “I use chatbots for world-building inspiration—creating characters and scenarios for my Dungeons and Dragons campaigns.”
Personal issues and mental health (14.17%) are notable topics, with users valuing the non-judgmental nature of chatbots: “I can talk about my problems, and it’s like having a private conversation with no fear of being criticized.” One user shared how a companion chatbot becomes an option for venting when people are not an option, stating, “I struggle with social anxiety, and when I can’t talk to anyone else, I use a chatbot to vent my feelings.”
Users also discuss past events (12.60%) and future plans (11.70%), using chatbots as tools for reflection and planning. One user explained, “I talk to the chatbot about my future goals and plans because it helps me organize my thoughts.”
Interpersonal issues and drama (10.96%) are explored with chatbots, allowing users to gain perspective on social conflicts: “When I have drama with people in my life, I talk to the chatbot to sort through my feelings.”
@@
Outside of these common themes, users discuss a wide range of niche topics with companion chatbots, from hobby-related questions to niche curiosities. One individual mentioned, “I use the chatbot to ask about random things like historical facts or recommendations for books. It’s like having a personal assistant for any question I have.” Others use companion chatbots as tools for specific tasks, such as recipe ideas.
@@
This diverse range of motivations and conversation topics highlights the versatility of companion chatbots in meeting users’ needs for entertainment, emotional support, and intellectual stimulation.
4.2 Correlations in psychological, social, and companion chatbot usage characteristics
In our analysis of factors that may influence chatbot usage, we calculated a Spearman correlation matrix with a Benjamini-Hoch correction to decrease the false discovery rate. The significant correlations after this adjustment are shown in Figure 8. While we note that these findings represent correlations, not causation, several interesting insights emerged from the correlation analysis regarding chatbot usage, loneliness, and related factors. A list of significant correlations and the corrected p-values with our key variables – usage (frequency and session length) and loneliness (as measured by ULS-8) – can be seen in Appendix A.
@@ 4.2 Correlations in psychological, social, and companion chatbot usage characteristics
For our prespecified analyses of usage vs. loneliness, we used uncorrected p-values. We found a significant if small correlation between the length of chatbot sessions and loneliness ($r=0.101,p=0.042$), but not between loneliness and the frequency of chatbot usage ($r=0.053,p=0.287$), as shown in Figure 9 and Figure 10. This suggests that feelings of loneliness are associated with how long people engage with chatbots but not how often.
For our exploratory analysis to discover variables correlated with usage and loneliness, we used the corrected p-values. The length and frequency of chatbot sessions showed moderate positive correlations with users’ perceptions of the chatbot. Specifically, longer sessions and higher frequency were associated with higher ratings of the state empathy towards a chatbot ($r>0.3,p<0.001$ for both) and attraction towards a chatbot ($r>0.18,p<0.001$ for all except frequency vs. task attraction, which had $r=0.165,p=0.002$). This suggests that users are more likely to use chatbots more when they feel it is empathetic, socially appealing, and relatable.
In terms of personality traits, loneliness was negatively correlated with extraversion ($r=-0.188,p<0.001$) and conscientiousness ($r=-0.335,p<0.001$), and positively correlated with neuroticism ($r=0.526,p<0.001$) and agreeableness ($r=0.333,p<0.001$). These findings suggest that individuals who are more lonely tend to be notably less extraverted and conscientious, and more neurotic and agreeable. Interestingly, these results match Buecker et al.’s [Buecker et al.(2020)] meta-analysis in all traits except agreeableness.
System usability, as measured by the System Usability Scale (SUS), showed a moderate positive correlation with chatbot session length ($r=0.196,p<0.001$) and usage frequency ($r=0.246,p<0.001$). This suggests that users who find the system more usable tend to engage in longer sessions but may not necessarily use the chatbot more frequently.
The problematic use of chatbots, measured by the GPIUS-2 scale, was moderately correlated with session length ($r=0.303,p<0.001$) and frequency ($r=0.337,p<0.001$). This indicates, unsurprisingly, that users who spend more time in chatbot sessions experience more problematic use.
These findings suggest that loneliness is associated with longer session length with companion chatbots, though many other variables may play a part in determining these variables and mediating their relationship. It is important to note that these findings represent correlations, not causation.
Figure 8: Correlation matrix of psychological, social, and usage characteristics.
Figure 9: Left: Chatbot Usage Frequency and Loneliness measured by ULS-8 scale. Right: Chatbot Session Length and Loneliness measured by ULS-8 scale.
@@
Figure 10: Left: Chatbot Usage Frequency and Problematic Use (GPIUS-2). Right: Chatbot Session Length and Problematic Use (GPIUS-2).
4.3 Multiple Regression Model

### v2

@@ 3.1.2 Custom Survey Items
In addition to established scales, we developed custom items to capture specific aspects of chatbot usage. Each item except usage characteristics required participants to elaborate on their choices and share thoughts not captured in the multiple-choice options as a free response, as our provided options were not expected to be exhaustive.
  • Usage Characteristics: Multiple-choice questions for frequency of chatbot use (5 options ranging from daily to once a month or less plus options indicating variation) and average session length (6 options: <15 min, 15-30 min, 10 min-1h, 1h-2h, 2h-4h, 4h+). Broad options were chosen to reduce the risk of inaccuracies from recall bias.
  • Motivations and Continued Use: Multiple-choice questions with required free-text elaboration on why participants began using chatbots and why they continue to use them; the options can be seen in Figure 5 (left graphs). These options were chosen through discussion with multiple researchers and based on uses and gratifications theory [Katz et al.(1973), Tanta et al.(2014)].
  • Conversation Topics: Questions about the topics participants discuss with chatbots, with the options chosen from discussions with multiple researchers (see top right graph of Figure 5).
  • Human vs. Chatbot Preference: Several items for participants’ preferences for human or chatbot interaction in various contexts, with the above conversation topics presented as options in a matrix question.
  • Perceived Impact on Relationships: Six Likert scales assessing participants’ perceptions of how chatbot use has affected their relationships with people by leading to (1) less interaction, (2) changes in conversation topics, and (3) greater ease, (4) greater enjoyment, (5) less engagement, and (6) less enjoyment in talking to people. These were selected through discussion with multiple researchers to roughly gauge participants’ perceived relationship changes.
@@ 3.2 Correlation and Multiple Regression
Data processing and statistical analyses were performed using Python, with the statsmodels library for regression analyses and scikit-learn for data preprocessing. We used a significance level of $\alpha=0.05$ for all statistical tests, unless otherwise specified. Usage frequency and session length were converted to linear numerical scores, with multiple responses averaged. All numerical variables were standardized using StandardScaler to improve interpretability and reduce potential multicollinearity [Aiken and West(1991)].
We computed a Spearman correlation matrix [Cohen et al.(2003)] with Benjamini-Hochberg correction for multiple comparisons[Benjamini and Hochberg(1995)]. For the regression analysis, we focused on session length as the independent variable, as it showed a significant correlation with loneliness where frequency did not.
@@
Potential mediators were identified using the corrected p-values (p < 0.05). Mediation analysis followed Baron and Kenny’s [Baron and Kenny(1986)] approach, with Sobel tests [Sobel(1982)] assessing mediation significance. We tested moderation effects by creating interaction terms and examining their significance in regression models [Aiken and West(1991)]. To address potential multicollinearity, we calculated Variance Inflation Factors (VIF), considering variables with VIF > 5 for elimination [Hair et al.(2010), O’brien(2007)]. After removing the state empathy subscales, which were redundant with the full scale, all predictors were below the threshold.
@@
To build our regression model, we started with all identified mediators, moderators, and confounders [Greenland et al.(1999), VanderWeele and Shpitser(2013)], including socialization and neuroticism on theoretical grounds. While retaining age and sex as control variables, we used backward elimination to remove predictors with high p-values (generally p > 0.1) to make a more parsimonious and interpretable model [Cohen et al.(2003)]. We then explored adding interaction terms one at a time, retaining significant interactions. Categorical variables (area type, education, relationship status, political party, race, employment status) were tested using ANOVA but were not retained due to non-significance.
@@
To examine the potential mediators of our final model, we conducted a parallel multiple mediation analysis. This analysis was performed using a bootstrapping approach with 5,000 resamples, as recommended by [Preacher and Hayes(2008)]. The analysis controlled for several covariates, including age, sex, and personality traits (BFI). Confidence intervals (95%) for the indirect effects were calculated using the percentile method.
@@ 3.3 Cluster Analysis
We employed K-means clustering analysis to identify distinct user profiles of chatbot users based on their psychosocial characteristics. These profiles were identified based on a set of features with significant correlations (p > 0.05, corr > 0.3) to key variables, as well as features relevant to chatbot usage including usage frequency and measures of empathy and attraction with a chatbot; the features were reduced to 10 principal components. After some exploration, we used Chatbot Session Length, Frequency, Loneliness (ULS-8), and Problematic Use of Chatbot (GPIUS2) for our key variables. The optimal number of clusters was determined using a combination of quantitative metrics and qualitative interpretation:
@@
The final clustering was performed using the K-means algorithm with k=7, implemented in Python using the scikit-learn library [Pedregosa et al.(2012)]. The resulting clusters were then characterized based on their centroids and distribution of key variables. To validate our clustering solution, we employed a multi-faceted approach:
@@
    After confirming significant differences with Kruskal-Wallis tests, we performed Dunn’s post-hoc test with Bonferroni correction for multiple comparisons [Dunn(1964)] to identify specific significant differences between clusters on our key variables (session length, loneliness, problematic use, and frequency).
  3. 3.
@@ 3.4 Interpretation of Quantitative Results and Free-Text Responses
Second, we examined the free-text responses associated with multiple-choice items to complement our quantitative findings. We conducted a thematic analysis [Braun and Clarke(2006)] of three free-response items (participants’ reasons for initial interest, motivations for continuing usage, and topics of conversation) using a stratified random sample; 15 participants were randomly selected from each cluster while maintaining equal representation across clusters (total N = 105). A researcher read through the sample of free responses, developed a codebook, and extracted overall themes for each examined item. The researcher developed initial codes by identifying recurring patterns, then grouped these codes into broader themes that captured key aspects of users’ reasons for engaging with companion chatbots.
@@ 4 Results
4.1 RQ1: What motivates the use of companion chatbots, and what do people use them for?
Figure 4: Percentage of total responses for multiple-choice questions regarding chatbot usage patterns among participants. Top left: chatbot session length. Top right: chatbot usage frequency. Bottom left: chatbot usage timeframe. Bottom right: most used chatbot platforms.
Figure 5: Percentage of participant responses for multiple-choice questions regarding their use of companion chatbots. Top left: reasons for using companion chatbots. Top right: conversation topics with chatbots. Bottom left: motivations for continued use. Bottom right: preferred devices for companion chatbot use.
Figure 6: Codes and themes for initial and continued usage of companion chatbots from thematic analysis of a stratified random sample of participants (15 per cluster, total N = 105).
4.1.1 Motivations for initial interest and continued use of companion chatbots
Analysis of both quantitative and qualitative data revealed diverse motivations for chatbot engagement, with several key themes emerging across both initial adoption and continued use. Descriptions of these themes can be found in Figure 6.
Technology exploration emerged as the predominant motivation, manifesting quantitatively in curiosity about AI and chatbots (30.89% initial use, 20.91% continued use). Users expressed interest in testing the technology’s capabilities and limitations: “I wanted to see if it could actually emulate a real person.”
Recreation represented another major theme, encompassing both entertainment (26.35% initial, 23.69% continued) and creative activities (17.81% initial use). Users enjoyed role-playing with characters and leveraging chatbots for creative projects: “I use chatbots to practice conversations for my D&D game, creating NPC personalities and dialogue.”
The practical utility theme emerged particularly in continued use, with users valuing chatbots for task assistance and information seeking. As one user explained: “I usually use chatbots to help with ideas like dinner ideas or information I may want to find out without using a typical search engine.” This theme aligns closest with the quantitative findings showing passing time and reducing stress (12.88% continued use) as a notable factor, though there were no directly-related multiple-choice options originally.
Another theme not explicitly reflected in the initial multiple-choice options was the safe space theme, which emerged to encompass the value of chatbots as non-judgmental environments for self-expression an emotional support. One participant noted, “I sometime feel lonely and just want to be left alone, during this time I like chatting with my AI companion because I feel safe and won’t not [sic] be judged for the inadequate decisions I have made.”
Social motivations, while less prevalent, are significant for some users. These motivations encompassed themes companionship, self-improvement, and lifestyle integration. Initial use showed loneliness (12.24%), social skill development (7.24%), and relationship seeking (3.53%) as drivers, while continued use revealed ongoing needs for companionship (9.86%), emotional experiences (6.52%) and self-discovery (6.12%). Some users explicitly sought social improvement: ‘I’m autistic, so I’m not the best at social interaction, and I found that practicing with chatbots helped.” The lifestyle integration theme emerged specifically in continued use, with some users incorporating chatbots into their daily routines: “It has become part of my life and I want to continue to learn and evolve with new technology.” This suggests that for some users, chatbots transition from novelty to utility and integration over time.
Our thematic analysis enriches our understanding beyond pure statistics, revealing how different motivations interweave and evolve throughout users’ engagement with companion chatbots. While initial interest often stems from curiosity and entertainment, continued use appears more deeply rooted in practical and emotional benefits.
Figure 7: Codes and themes for topics of conversation with companion chatbots from thematic analysis of a stratified random sample of participants (15 per cluster, total N = 105).
@@ 4.1.2 Topics of conversation with companion chatbots
Analysis of both structured responses and free-text data revealed five major themes in how users engage with companion chatbots (see Figure 7).
Casual exchange, including casual conversations (26.28%) and entertainment (21.66%), represents the most common form of interaction. Participants used the chatbot to review past events (12.60%), such as talking about their day: “I often chat with my AI about random things like my day, and it feels like I’m having a normal conversation without the pressure of talking to a real person.”
Emotional disclosure emerged as a notable theme, encompassing mental health discussions and personal problems. This is reflected in the quantitative findings of personal issues and mental health (14.17%) and interpersonal issues and drama (10.96%) as notable topics. One participant responded, “I can talk about my problems, and it’s like having a private conversation with no fear of being criticized.”
Though less prevalent in the free responses, the theme of creative development manifested through topics of writing, roleplay, and creative ideation, with users leveraging chatbots as collaborators: “I typically ask for ideas for my marketing art! I do a lot of graphic design for a lot of different people and sometime run low on creative ideas.”
@@
Knowledge seeking involved users turning to chatbots for information and advice, including discussion of future plans (11.70%). One individual mentioned, “I use the chatbot to ask about random things like historical facts or recommendations for books.” Others used companion chatbots as tools for specific tasks, such as recipe ideas.
A rare theme emerging from the free responses was intimate exchange, involving romantic and sexual interactions: “[Some]times I just want to feel pleasure and the interaction is very sexual. Other times I want to feel love and connection and it’s more romantic.”
@@
We note that the themes identified in free responses sometimes diverge from the structured response options. This diverse range of conversation topics highlights the versatility of companion chatbots in meeting users’ needs for entertainment, emotional support, and intellectual stimulation.
4.2 RQ2: What is the relationship between companion chatbot usage and loneliness, and what factors mediate or moderate this relationship?
4.2.1 Correlations in psychological, social, and companion chatbot usage characteristics
@@
For our prespecified analyses of usage vs. loneliness, we found a small but significant correlation between chatbot session length and loneliness ($r=0.101,p=0.042$), but not between loneliness and usage frequency ($r=0.053,p=0.287$), as shown in Figure 9.
Several other notable correlations emerged in our exploratory analysis using Benjamini-Hochberg corrected p-values. Chatbot usage (both length and frequency) showed moderate positive correlations with state empathy towards chatbots ($r>0.3,p<0.001$) and various forms of attraction towards chatbots ($r>0.18,p<0.001$). Loneliness demonstrated significant correlations with personality traits of extraversion, ($r=-0.188,p<0.001$), conscientiousness ($r=-0.335,p<0.001$), agreeableness ($r=-0.333,p<0.001$), and neuroticism ($r=0.526,p<0.001$), aligning with Buecker et al.’s [Buecker et al.(2020)] meta-analysis. Additionally, problematic use of chatbots (GPIUS2) showed moderate correlations with both session length ($r=0.303,p<0.001$) and frequency ($r=0.337,p<0.001$). See Figure 8 for a correlation matrix of our continuous variables.
While these correlations provide initial insights into the relationships between variables, they represent associations rather than causation. A more detailed examination of these relationships through multiple regression analysis follows.
Figure 8: Correlation matrix of psychological, social, and usage characteristics. Variables are sorted into general categories, and names have been altered for readability. The original names can be seen in Figure 18 of Appendix A.
@@
Figure 9: Plots of loneliness (ULS-8) and problematic usage (GPIUS2) vs. usage (frequency and session length). Top left: loneliness vs. chatbot usage frequency. Top right: loneliness vs. chatbot session length. Bottom left: problematic use vs. chatbot usage frequency. Bottom right: problematic use vs. chatbot session length..
Figure 10: The final model from our exploratory multiple regression, depicting the relationships between chatbot session length, loneliness, and related factors.
4.2.2 Multiple Regression Model